Q11:- At 400 K, the half-life for the decomposition of a sample of a gaseous compound initially at 55.5 kPa was 340 s. When the pressure was 28.9 kPa, the half-life was 178 s. Determine the order of the reaction.
1
Expert's answer
2020-12-29T02:42:52-0500

Let's test n=0:

"k_1=(p_0)_1\/(2*t_{1\/2})=55500\/(2*340)=81.6"

"k_2=(p_0)_2\/(2*t_{1\/2})=28900\/(2*178)=81.1"

Constants are almost equal, so we have 0-order reaction.